# Fully Spiking Neural Network for Legged Robots

### v2

@@ Abstract
In recent years, legged robots based on deep reinforcement learning have made remarkable progress. Quadruped robots have demonstrated the ability to complete challenging tasks in complex environments and have been deployed in real-world scenarios to assist humans. Simultaneously, bipedal and humanoid robots have achieved breakthroughs in various demanding tasks. Current reinforcement learning methods can utilize diverse robot bodies and historical information to perform actions. However, prior research has not emphasized the speed and energy consumption of network inference, as well as the biological significance of the neural networks themselves. Most of the networks employed are traditional artificial neural networks that utilize multilayer perceptrons (MLP). In this paper, we successfully apply a novel Spiking Neural Network (SNN) to process legged robots, achieving outstanding results across a range of simulated terrains. SNN holds a natural advantage over traditional neural networks in terms of inference speed and energy consumption, and their pulse-form processing of body perception signals offers improved biological interpretability. Applying more biomimetic neural networks to legged robots can further reduce the heat dissipation and structural burden caused by the high power consumption of neural networks. To the best of our knowledge, this is the first work to implement SNN in legged robots.
@@
[…]
I Introduction
The increasing adoption of mobile robots, which are equipped with continuous high-dimensional observations and action space, to tackle a wide range of intricate tasks in real-world situations emphasizes the critical importance of exceptional control algorithms. Currently, the limited onboard energy resources of most robots pose a significant challenge, as this constraint hinders their ability to operate continuously and cost-effectively. Consequently, there is an immediate demand for developing energy-efficient solutions for the seamless control of these autonomous machines. Deep reinforcement learning (DRL) employs deep neural networks (DNNs) as potent function approximators for learning optimal control strategies for intricate tasks [1, 2], through directly mapping the original state space to the action space [3, 4]. Nonetheless, the remarkable performance of DRL frequently comes at the expense of substantial energy consumption and slower execution speeds, making them unsuitable for various applications. Additionally, the execution speed of control strategies employing DNNs tends to be slower in comparison to the operational speed of the motion units. This discrepancy often results in a step-like behavior in the control signals, causing negative impacts on the performance of the system. Moreover, the reliance of current Deep Neural Networks (DNNs) on high-power GPUs for computation significantly impacts the structural design of robots, necessitating the inclusion of substantial active cooling systems.
@@ I Introduction
Spiking neural networks (SNNs), also referred to as third-generation neural networks, present a promising alternative for energy-efficient and high-speed deep networks. These emerging SNNs operate based on the principles of neuromorphic computing, wherein the integration of memory and computation is seamless, and neurons engage in asynchronous, event-based communication and computation [5]. The biological plausibility, the significant increase in energy efficiency (particularly when deployed on neuromorphic chips [6]), high-speed processing and real-time capability for high-dimensional data (especially from asynchronous sensors like event-based cameras [7]) contribute to the advantages that SNNs possess over ANNs in specific applications. These advantages render the utilization of SNNs not only feasible but also highly advantageous in lieu of ANNs for performing effective calculations. A mounting body of research illustrates that SNNs can function as energy-efficient and high-speed solutions for effectively managing robot control in scenarios where there are limitations on onboard energy resources [8, 9, 10]. To address the limitations of SNNs in tackling high-dimensional control problems, a natural approach involves combining the energy efficiency of SNNs with the optimality of DRL, which has proven effective in various control tasks [11]. Due to the role of rewards as training guides in reinforcement learning (RL), some studies utilize a three-factor learning rule [12] to implement reward learning. Although these rules exhibit strong performance in low-dimensional tasks, they often struggle to handle complex problems, and the optimization process becomes challenging in the absence of a global loss function [13]. Recently, [14] proposed a strategy gradient-based algorithm to train an SNN for learning random strategies. However, this algorithm is designed for discrete action spaces, and its practical applications are somewhat limited when tackling high-dimensional continuous control problems.
@@
The recent conceptualization of the brain’s topology and computational principles has ignited advancements in SNNs, exhibiting both human-like behavior[15] and superior performance[16]. A pivotal attribute associated with effective computation in the brain is the employment of neuronal populations for encoding and representing information, encompassing the entire spectrum from sensory input to output signals. In this scenario, each neuron within a population has a receptive field that captures a specific segment of the encoded signal[17]. Notably, initial investigations into this group coding scheme have shown its enhanced capability to represent stimuli[18], contributing to recent triumphs in training SNNs for complex, high-dimensional supervised learning tasks[19, 20]. The efficacy of population coding presents a promising pathway for the advancement of efficient SNNs that leverage population coding. These networks possess the potential to acquire optimal solutions for complex high-dimensional continuous control tasks, paving the way for significant progress in this field.
The main contributions of this paper can be summarized as follows:
@@
  • For the first time, we have implemented SNNs on a policy network in various legged robots simulated in Isaac Gym[21], enabling the encoding of each dimension of the observation and action space within a single population of neurons using a learnable acceptance domain. Furthermore, we have successfully integrated this method and achieved successful training outcomes using techniques such as imitation learning and trajectory history.
@@
  • Our approach presents a considerable advantage over conventional ANNs in terms of energy efficiency. This advantage holds substantial significance for enhancing the structural integrity and reducing the costs associated with robot development.
@@
  • The research affirms the exceptional performance demonstrated by SNNs in the domain of high-frequency robot control, coupled with their significant edge over ANNs in attenuating signal noise, which enhances their robustness in practical situations.
II Related work
II-A Reinforcement Learning for Legged Robotics Locomotion
In recent years, there has been a rapid development of reinforcement learning techniques in the realm of achieving fast, stable, and adaptive locomotion for legged robots. [22] bridges the gap between reinforcement learning simulation and actual reality by modeling the actuator through neural networks. [23] utilizes the teacher-student training framework to integrate environmental parameters with preconceptions which enables quadruped robots to adapt to complex terrains. [24] introduces a large number of auxiliary rewards and gait parameter control to implement multiple gaits for a single policy. This work brings a lot of inspiration to the design of quadruped rewards. [25] combines reinforcement learning with computer vision to complete simultaneous locomotion and manipulation. Unlike traditional reinforcement learning, Generative adversarial imitation learning (GAIL) is an approach for imitating behaviors from reference datasets by integrating a generative adversarial network. Based on the GAIL, Adversarial Motion Priors (AMP) [26] combines the task reward and imitation reward to make the agent complete the tasks based on the action being similar to the reference dataset. For learning unlabeled references dataset, [27] uses a skill discriminator additionally and enables quadruped to distinguish and master multiple gaits, and perform backflips not found in the dataset. [28] combines RMA and AMP, allowing the quadruped with the ability to traverse challenging terrains rapidly. However, the above methods are realized by ANNs. Therefore, it cannot take into account the high-frequency and energy-saving advantages of SNNs.
@@ II-A Reinforcement Learning for Legged Robotics Locomotion
Figure 2: The observations are initially encoded by the encoder as $n$ independent distributions that are uniformly distributed over the observation range. After encoding, the population processes the distributions, resulting in spike generation. The neurons in the input populations encode each observation dimension and drive a multi-layered, fully connected SNN. During forward timesteps in PopSAN, the activities of each output population are decoded to determine the corresponding action dimension. This implies that the neural network receives observations, processes them using the SNN, and decodes the resulting activities to determine the appropriate action for the specific situation.
II-B Spiking Neuron Networks
Recently, many works have grown up around introducing SNNs into RL algorithms[29, 30, 31, 32, 12]. SNNs offer advantages in processing temporal information, energy efficiency, event-driven processing, robustness to noise, plasticity, and biological plausibility. Some methods get trained ANNs to be converted into corresponding SNNs with few accuracy loss by matching the firing rates of spiking neurons with the graded activation of analog neurons[33].
Yet following the surrogate gradient method[34], the spike-based backpropagation (BP) algorithm has quickly become the mainstream solution for training multi-layer SNNs[35]. As shown in several open-source frameworks of SNNs[36, 37], the membrane voltage of non-spiking neurons is feasible to represent a continuous value in a spike-based BP method.
III SNNs based Locomotion in Isaacgym
We opted to evaluate the efficacy of our algorithm by training and testing it using Isaac Gym[21]. Isaac Gym is a simulation platform specifically created for robotics applications. The software provides a realistic physics simulation, specialized support for legged robots, integration with NVIDIA technologies, customizable features, and a vibrant community.
III-A SNN based Policy Network
Algorithm 1 Forward propagation through PopSAN
  Randomly initialize weight matrices $\bm{W}$ and biases $\bm{b}$ for each SNN layer;
  Initialize encoding means $\bm{\mu}$ and standard deviations $\bm{\sigma}$ for all input populations;
  Randomly initialize decoding weight vectors $\bm{W}_{d}$ and bias $b_{d}$ for each action dimension;
  N-dimensional observation, $\bm{s}$;
  Spikes from input populations generated by encoder module: $\bm{X}$ = Encoder($\bm{s}$, $\bm{\mu}$, $\bm{\sigma}$);
  for $t=1,...,T$ do
     Spikes at timestep t: $\bm{o}^{(t)(0)}=\bm{X}^{(t)}$;
     for $k=1,...,K$ do
        Update LIF neurons in layer $k$ at timestep $t$ based on spikes from layer $k-1$:
        $\bm{c}^{(t)(k)}=d_{c}\cdot\bm{c}^{(t-1)(k)}+\bm{W}^{(k)}\bm{o}^{(t)(k-1)}+\bm{b}^{(k)}$;
        $\bm{v}^{(t)(k)}=d_{c}\cdot\bm{v}^{(t-1)(k)}\cdot(1-\bm{o}^{(t-1)(k)})+\bm{c}^{(t)(k)}$;
        $\bm{o}^{(t)(k)}=Threshold(\bm{v}^{(t)(k)})$;
     end for
  end for
  $M$-dimensional action $\bm{a}$ generated by decoder module:
  Sum up the spikes of output populations: $\bm{sc}=\sum_{t=1}^{T}\bm{o}^{(t)(K)}$;
  for $i=1,...,M$ do
     Compute firing rates of the $i^{th}$ output population: $\bm{fr}^{(i)}=\bm{sc}^{(i)}/T$;
     Compute $i^{th}$ dimension of action: $\alpha^{i}={\bm{W}_{d}}^{(i)}\cdot\bm{fr}^{(i)}+{b_{d}}^{(i)}$;
  end for
@@ III-A SNN based Policy Network
We employ a population-coded spiking actor-network (PopSAN)[38] that is trained in tandem with a deep critic network using the DRL algorithms (including RMA and AMP). During training, the PopSAN generated an action $\alpha$in $\mathbb{R}^{N}$ for a given observation, $s$, and the deep critic network predicted the associated state value $V$($s$) or action-value $Q$($s$, $\alpha$), which in turn optimized the PopSAN, in accordance with a chosen DRL method (Fig. 2). Within the PopSAN architecture, the encoder module is responsible for encoding individual dimensions of the observation by mapping them to the activity of distinct neuron populations. During forward propagation, the input populations activate a multi-layer fully-connected SNN. The SNN then produces activity patterns within the output populations. At the end of each set of $T$ timesteps, these patterns of activity are decoded in order to ascertain the associated action dimensions. (as outlined in Algorithm 1).
@@
The current-based leaky-integrate-and-fire (LIF) model of a spiking neuron is employed in constructing the SNN. This model is utilized in constructing the SNN architecture. The dynamics of the LIF neurons are governed by a two-step model, as elaborated in Algorithm 1: i) the integration of presynaptic spikes $o$ into current $c$; and ii) the integration of current $c$ into membrane voltage $v$; $d_{c}$ and $d_{v}$ represent the current and voltage decay factors, respectively. In this particular implementation, a neuron fires a spike when its membrane potential surpasses a predetermined threshold. The hard-reset model was implemented, in which the membrane potential is promptly reset to the resting potential following a spike. The resultant spikes are transmitted to the post-synaptic neurons during the same inference timestep, under the assumption of zero propagation delay. This approach facilitates efficient and synchronized information transmission within the SNN.
@@
Figure 3: RMA consists of two subsystems: the base policy $\pi$ and the adaptation module $\phi$. The RMA training consists of two phases. Training the Base Policy (Phase 1): In the initial phase, the base policy $\pi$ is trained using PopSAN. The system takes the current state $x_{t}$, the previous action $\alpha_{t-1}$, and the environmental factors $e_{t}$ as input. These environmental factors are encoded into a latent extrinsics vector $z_{t}$ using the environmental factor encoder $\mu$. Training the Adaptation Module (Phase 2):In the second phase, the adaptation module $\phi$ is trained to predict the extrinsics $\widehat{z_{t}}$ using past states and actions. This training utilizes supervised learning with on-policy data. The adaptation module learns to capture the relationship between the state-action history and the corresponding extrinsics. By training the base policy and the adaptation module in these two phases, RMA can learn and adapt to the environment more effectively.
@@
Next, we combined the advantages of SNN with the advanced RMA algorithm. Figure 3 shows that the RMA system consists of two interconnected subsystems: the base policy $\pi$ and the adaptation module $\phi$. These subsystems work together to support continuous adaptation in various environmental conditions, ensuring seamless online operations. The base policy is trained using reinforcement learning in simulation, leveraging privileged information about the environment configuration $e_{t}$, such as friction, payload, and other factors. By utilizing the vector $e_{t}$, the base policy can adapt effectively to the unique characteristics of the environment.
Unfortunately, implementing this policy directly is not feasible in reality because $e_{t}$ is unavailable. To address this challenge, we must estimate the extrinsics during runtime, a task performed by the adaptation module $\phi$. The actual movement of robot joints deviates from the intended movement due to extrinsic influences. Instead of depending on privileged information, we can use the agent’s recent state history to estimate the extrinsics vector. The purpose of $\phi$ is to estimate the extrinsics vector $z_{t}$ based solely on the recent state and action history of the robot, without direct access to $e_{t}$.
@@
In addition, we have successfully combined SNN with AMP and achieved similar performance to ANN on legged robots. Figure 4 provides a schematic overview of the system. The motion dataset $M$ consists of a collection of reference motions, where each motion $m^{i}$ $=$ $\widehat{q}_{t}^{i}$ is represented as a sequence of poses $\widehat{q}_{t}^{i}$. The motion clips can be obtained from a variety of sources, including motion capture (mocap) recordings of real-life actors or artist-authored keyframe animations. The simulated robot’s movement is governed by a policy $\pi(\alpha_{t}|s_{t},g)$ that links the character’s state $s_{t}$ and a given goal $g$ to a distribution of actions $\alpha_{t}$. The policy generates actions that determine the desired target positions for proportional-derivative (PD) controllers at each joint of the robot. The controllers then generate control forces to propel the robot’s motion according to specified target positions. The goal $g$ defines a task reward function $r_{t}^{G}=r^{G}(s_{t},\alpha_{t},s_{t+1},g)$ that outlines the high-level objectives the robot needs to achieve. The style objective $r_{t}^{S}=r^{S}(s_{t},s_{t+1})$ is determined by an adversarial discriminator, which provides an a priori estimate of the naturalness or style of a motion, independent of the task at hand. By doing this, the style objective encourages the policy to produce movements that closely mirror the behaviors seen in the dataset.
@@
III-B Training of Legged Robot using SNN on Isaacgym

### v3

@@ Abstract
Recent advancements in legged robots using deep reinforcement learning have led to significant progress. Quadruped robots can perform complex tasks in challenging environments, while bipedal and humanoid robots have also achieved breakthroughs. Current reinforcement learning methods leverage diverse robot bodies and historical information to perform actions, but previous research has not emphasized the speed and energy consumption of network inference and the biological significance of neural networks. Most networks are traditional artificial neural networks that utilize multilayer perceptrons (MLP). This paper presents a novel Spiking Neural Network (SNN) for legged robots, showing exceptional performance in various simulated terrains. SNNs provide natural advantages in inference speed and energy consumption, and their pulse-form processing enhances biological interpretability. This study presents a highly efficient SNN for legged robots that can be seamless integrated into other learning models.
I Introduction
The increasing adoption of mobile robots with continuous high-dimensional observations and action spaces necessitates advanced control algorithms for complex real-world tasks. Currently, the limited onboard energy resources hinder continuous and cost-effective operation, creating an urgent need for energy-efficient solutions for the seamless control of these robots. Deep reinforcement learning (DRL) employs deep neural networks (DNNs) as potent function approximators for learning optimal control strategies for intricate tasks [1, 2], through mapping the original state space to the action space [3, 4]. [5] differs from traditional reinforcement learning by imitating behaviors from reference datasets through a generative adversarial network. Adversarial Motion Priors (AMP) [6] enhances [5] by combining task and imitation rewards, enabling agents to mimic actions from reference datasets. To learn from unlabeled references dataset, [7] employs a skill discriminator, allowing quadrupeds to master various gaits and perform backflips. Furthermore, [8] integrates Rapid Motor Adaptation (RMA) with AMP, improving quadrupeds’ ability to traverse challenging terrains rapidly. While DRL delivers impressive performance, it often incurs high energy consumption and slower execution speeds. DNN-based control strategies generally operate slower than motion units, causing step-like control signals that hinder performance.
@@ I Introduction
Spiking neural networks, or third-generation neural networks, provide an energy-efficient and high-speed alternative for deep learning by utilizing neuromorphic computing principles [9]. The biological plausibility, the significant increase in energy efficiency (particularly when deployed on neuromorphic chips [10]), high-speed processing and real-time capability for high-dimensional data (especially from asynchronous sensors like event-based cameras [11]) contribute to the advantages that SNNs possess over ANNs in specific applications. Recently, many works have grown up around introducing SNNs into RL algorithms[12, 13, 14, 15, 16]. Research shows that SNNs are energy-efficient and high-speed solutions for robot control in scenarios with limited onboard energy resources [17, 18, 19]. To address the limitations of SNNs in high-dimensional control problems, combining their energy efficiency with the optimality of DRL offers a promising solution, as DRL has proven effective in various control tasks [20]. Rewards act as training guides in DRL, some studies utilize a three-factor learning rule [16]. While effective in low-dimensional tasks, these rules struggle with complex problems, complicating optimization without a global loss function [21]. Recently, [22] proposed a strategy gradient-based algorithm for training a SNN to learn random strategies, but it is limited to discrete action spaces, hindering its use in high-dimensional continuous control problems.
@@
The recent conceptualization of the brain’s topology and computational principles has ignited advancements in SNNs, exhibiting both human-like behavior[23] and superior performance[24]. A key feature of brain computation is the use of neuronal populations to encode information, from sensory input to output signals. Each neuron has a receptive field that captures a specific segment of the signal [25]. Notably, initial investigations into this group coding scheme have shown its enhanced capability to represent stimuli[26], contributing to triumphs in training SNNs for complex, high-dimensional supervised learning tasks[27, 28]. The main contributions of this paper can be summarized as follows:
@@
  • For the first time, we have implemented a lightweight population coded SNNs on a policy network in various legged robots simulated in Isaac Gym[29] using a multi-stage training method. We also integrated this method with imitation learning and trajectory history, achieving effective training outcomes.
@@
  • Our approach presents a considerable advantage over ANNs in terms of energy efficiency. This advantage holds substantial significance for enhancing the structural integrity and reducing the costs associated with robot development.
@@
  • The research affirms the exceptional performance demonstrated by SNNs in high-frequency robot control, coupled with their significant edge over ANNs in attenuating signal noise, which enhances their robustness in practical situations.
@@
Figure 2: The observations are initially encoded by the encoder as $n$ independent distributions that are uniformly distributed over the observation range. After encoding, the population processes the distributions, resulting in spike generation. The neurons in the input populations encode each observation dimension and drive a multi-layered, fully connected SNN. During forward timesteps in PopSAN, the activities of each output population are decoded to determine the corresponding action dimension. The neural network receives observations, processes them using the SNN, and decodes the resulting activities to determine the appropriate action for the specific situation.
II Methods
II-A SNN based Policy Network
@@
We employ a population-coded spiking actor-network (PopSAN)[30] that is trained in tandem with a deep critic network using the DRL algorithms. During training, the PopSAN generated an action $\alpha$in $\mathbb{R}^{N}$ for a given observation, $s$, and the deep critic network predicted the associated state value $V$($s$) or action-value $Q$($s$, $\alpha$), which in turn optimized the PopSAN, in accordance with a chosen DRL method (Fig. 2). Within the PopSAN architecture, the encoder module encodes individual dimensions of the observation by mapping them to the activity of distinct neuron populations. During forward propagation, the input populations activate a multi-layer fully-connected SNN, producing activity patterns within the output populations. After each set of $T$ timesteps, these patterns of activity are decoded to ascertain the associated action dimensions.
@@
The current-based leaky-integrate-and-fire (LIF) model of a spiking neuron is employed in constructing the SNN. The dynamics of the LIF neurons are governed by a two-step model: i) the integration of presynaptic spikes $o$ into current $c$; and ii) the integration of current $c$ into membrane voltage $v$; $d_{c}$ and $d_{v}$ represent the current and voltage decay factors, respectively. In this implementation, a neuron fires a spike when its membrane potential surpasses a predetermined threshold. The hard-reset model was implemented, in which the membrane potential is promptly reset to the resting potential following a spike. Resultant spikes are transmitted to post-synaptic neurons during the same inference timestep, assuming zero propagation delay. This approach facilitates efficient and synchronized information transmission within the SNN.
II-B Temporal Shrinking
Next, inspired by [31], we process the encoded information from the encoder in i stages. At each subsequent stage, the time step can be reduced by an arbitrary scale. Assuming that the time step for the first stage is $T_{1}(T_{1}=n)$ and the scale for the second stage is $T_{2}(T_{2}=j)$, we utilize a learnable weight $W$in $\mathbb{R}^{T_{2}\times T_{1}}$ to perform the scale conversion, as illustrated below:
$I_{2}=O_{1}\odot Softmax(WPop_{mean}^{O_{1}})$ (1)
Where $I_{1}$in $\mathbb{R}^{T_{1}\times obs}$ represents the input of the first stage, which is at the same scale as the output $O_{1}$in $\mathbb{R}^{T_{1}\times obs}$ of the first stage, while the input of the second stage should be $I_{2}$in $\mathbb{R}^{T_{2}\times obs}$, and $Pop_{mean}^{O_{1}}$in $\mathbb{R}^{T_{1}\times 1}$ means the average of $O_{1}$ at each time point:
$Pop_{mean}^{O_{1}}=\frac{1}{obs}\sum_{p=0}^{obs}O_{1,p}$ (2)
The $obs$ in (2) refers to the number of elements observed at a single time step. Equation (1) utilize softmax function to ensure that the probabilities allocated at all time steps sum to 1, ensuring information integrity. This way, (2) can be seen as a guidance for (1) to learn how to allocate information from stage 1 to stage 2. The method enables infinite time step compression to a lightweight level at minimal cost, greatly affecting low-latency needs in robot gait control.
II-C Auxiliary Optimization
The use of surrogate gradient in training SNNs effectively tackles the non-differentiability of spikes, yet the discrepancy with the true gradient poses a limitation on SNN performance. Moreover, spikes face serious gradient vanishing/exploding problems [32]. To alleviate the aforementioned issues and ensure the effectiveness of temporal shrinking, we introduce auxiliary optimization. After each stage (except the final stage), $O_{i}$ is fed into the auxiliary classifier to obtain the stage loss in addition to the subsequent stage. The auxiliary classifier consists of an SNN classifier and a decoder module, where the time dimension of the SNN is equal to $T_{i}$ of the corresponding stage. The overall loss $L_{all}$ should be represented as a weighted sum of the losses from each stage, expressed as follows:
$L_{all}=\sum_{i}^{I}\lambda_{i}L_{i}(Y_{i},\mathbb{S}_{i}),\sum_{i}^{I}\lambda_{i}=1$ (3)
Where $\lambda_{i}$ is a manually set parameter representing the weight of each stage. $Y_{i}$ denotes the output of the auxiliary classifier in stage $i$, while $L_{i}$ represents the loss gained by interacting $Y_{i}$ with the environment $\mathbb{S}_{i}$ at this stage. By adjusting $\lambda_{i}$, we can emphasize either the global action features or the details of specific actions.
II-D Combination of Other Methods
@@
Figure 3: RMA consists of two subsystems: the base policy $\pi$ and the adaptation module $\phi$. The RMA training consists of two phases. Training the Base Policy (Phase 1): In the initial phase, the base policy $\pi$ is trained using PopSAN. The system takes the current state $x_{t}$, the previous action $\alpha_{t-1}$, and the environmental factors $e_{t}$ as input. These environmental factors are encoded into a latent extrinsics vector $z_{t}$ using the environmental factor encoder $\mu$. Training the Adaptation Module (Phase 2):In the second phase, the adaptation module $\phi$ is trained to predict the extrinsics $\widehat{z_{t}}$ using past states and actions. This training utilizes supervised learning with on-policy data. The adaptation module learns to capture the relationship between the state-action history and the corresponding extrinsics.
@@
To validate the generalizability of our method, we further combined the advantages of SNN with the advanced RMA algorithm. Figure 3 shows that the RMA system consists of two interconnected subsystems: the base policy $\pi$ and the adaptation module $\phi$. The base policy is trained using reinforcement learning in simulation, leveraging privileged information about the environment configuration $e_{t}$, such as friction, payload, and other factors. By utilizing the vector $e_{t}$, the base policy can adapt effectively to the unique characteristics of the environment. The purpose of $\phi$ is to estimate the extrinsics vector $z_{t}$ based solely on the recent state and action history of the robot, without direct access to $e_{t}$.
@@
In addition, we have successfully combined SNN with AMP and achieved similar performance to ANN on legged robots. Figure 4 provides a schematic overview of the system. The motion dataset $M$ consists of a collection of reference motions, where each motion $m^{i}$ $=$ $\widehat{q}_{t}^{i}$ is represented as a sequence of poses $\widehat{q}_{t}^{i}$. The simulated robot’s movement is governed by a policy $\pi(\alpha_{t}|s_{t},g)$ that links the character’s state $s_{t}$ and a given goal $g$ to a distribution of actions $\alpha_{t}$. The policy generates actions that determine the desired target positions for proportional-derivative (PD) controllers at each joint of the robot. The controllers then generate control forces to propel the robot’s motion according to specified target positions. The goal $g$ defines a task reward function $r_{t}^{G}=r^{G}(s_{t},\alpha_{t},s_{t+1},g)$ that outlines the high-level objectives the robot needs to achieve. The style objective $r_{t}^{S}=r^{S}(s_{t},s_{t+1})$ is determined by an adversarial discriminator, which provides an a priori estimate of the naturalness or style of a motion, independent of the task at hand. By doing this, the style objective encourages the policy to produce movements that closely mirror the behaviors seen in the dataset.
@@
II-E Training